# revtex4-2Repair the float

### v3

@@ 4.3 Patterns of the FT of the topological defects
We now analyze in detail the effects that the half-vortex and the double dislocation have on the induced CDW order. Recall that inside the core of the half-vortex the SC order parameter is mostly FF-type, since the amplitude of one of the two PDW order parameters must vanish at the location of the half-vortex, i.e. the origin (see section 3). Since the FF state breaks inversion symmetry in the $x$-direction, the corresponding CDW pattern inherits this broken symmetry. Also, the edge dislocations of the CDW order parameter breaks inversion symmetry in the $y$-direction. Note that, the parity operator, in the $x$-direction, changes the location of the branch cuts and the signs of the winding-numbers when it acts on the order parameters, which changes the sign of the dislocation-charge. As a result, the corresponding charge-density patterns are flipped on their head under this operation. Since the location of the branch cuts have no physically significant effects on the charge-density, the full-vortex is invariant under this operation.
The ordering wave vector $2\mathbf{Q}$ has many features made more apparent in the Fourier transforms of the LDOS which are shown in figures 6a-6b. Notice that the Fourier transforms of the LDOS for half-vortex and the double dislocation feature split peaks at the $2\mathbf{Q}$ ordering wave vector where the amplitude of the FT-LDOS is zero there. In contrast, in the case of the Abrikosov vortex, the Fourier Transform of the LDOS is just the transform of $\cos(2\mathbf{Q}x)$ which has a single peak at the ordering wave vector. Cuts of the FT-LDOS along $k_{y}=0$ are given in figures 6c and 6d to more clearly illustrates these split peaks.
@@
The split peaks are signature of the defects of the PDW phase associated with jumps in the phase $\theta_{-}$, defined in equation (2.8), across the core of the topological defect. For example, across the half-vortex the phase jumps by ±$\pm\pi/2$ since it winds by $\pi$ around the half-vortex. This implies that the $2Q$ Fourier component is equal to itself times $i$ across the core of the defect, suggesting that this Fourier component must be zero. We can explicitly verify this prediction by examining the FT of $\rho_{2\mathbf{Q}}(\mathbf{r})$. It is apparent that a nonzero CDW-winding-number is responsible for the vanishing of $\rho_{2\mathbf{Q}}(\mathbf{r})$ at the center of the defect. In other words, the phase shift that causes this destructive interference is a measurement of the Burgers vector which is the topological charge of the dislocation. Similar interference patterns are well known to exist in electron diffraction in crystals of semiconductors with dislocations.
@@
Similar split peaks in the Fourier transform of the tunneling LDOS were also predicted to exist at the PDW halo of an Abrikosov vortex of a superconductor in Refs. [16, 65] but their physical origin is very different. Indeed, in the case of the vortex halo there is a phase shift in the $\rho_{\mathbf{Q}}(\mathbf{r})$ (instead of $\rho_{2\mathbf{Q}}(\mathbf{r})$) Fourier component of the local charge density caused the Abrikosov vortex of the uniform component of the superconductor.
@@
Figure 8: Comparison of $\frac{dI}{dV}$ curves related to the electron tunneling DOS for a period $8$ PDW (LO state). The electron tunneling DOS is sensitive to the periodicity of the associated period $4$ CDW. The labeled $A-D$ are the tunneling DOS traces at the four inequivalent sites of the CDW. Each consecutive curve is offset by 1 for clarity. Notice that in the PDW traces particle-hole symmetry is only present at low bias (low energies).
@@
Figure 9: A set of $\frac{dI}{dV}$ curves for the half-vortex belonging to different lattice sites. The labels indicate the distance the lattice site in question is from the vortex core in fractions of the halo radius, $r_{0}=24a_{0}$. The top two curves are shifted up by .75.75.75.75 and $1.5$ units in respect to zero. Close to the center of the core the curves resemble a squeezed in FF state more so than an LO state, but as we move outward the coherence peaks grow, and the half-vortex behaves more like an LO state. In appendix E figure 14 additional curves corresponding to a half-vortex are provided which indicates a 4 lattice site periodicity, like the LO state, but there is also a shift in tunneling spectra associated with the jump in phase of $\theta_{+}$ across the vortex core. This, along with reduced coherence peaks and additional satellites, distinguished the half-vortex from the pure PDW even far outside the vortex core.
@@
Figure 10: The Bogoliubov Fermi surface, $A(\mathbf{k},\omega=0)$, for a PDW state with a half-vortex of radius $r_{0}=24a_{0}$ and a d-wave form factor. In (a) we plot the upper portion of the Bogoliubov Fermi surface, which indicates a redistribution of spectral weight in respect to the defect-free PDW (see appendix F for plots). In (b) we zoom in on these loops, which indicate inversion symmetry breaking. We also partition the Fermi surface of the half-vortex into (c) the particle-like portion and into (d) the hole-like portion. Notice the intensity of the spectral weight belonging to the Fermi arcs tend to be either particle-like or hole-like depending on the side you are on.
@@
Alternatively, when there is a winding-number in both PDW components, we can picture the phase jump as occurring in both winding-numbers independently. Recalling the form of the induced $2Q$-CDW from equation (2.6), we see that complex conjugation flips the phase-winding of one of the PDW components. Thus, when we have a vortex in both components $\Delta_{\pm\bm{Q}}$ with the same winding, i.e. the Abrikosov vortex, the phase differences cancel each other, and there is no net phase jump across the core of the vortex. Hence, in the case of the Abrikosov vortex there is not a split peak. Equivalently, the Abrikosov vortex does not have any dislocation charge associated with it to cause a split peak to exist. We can contrast this with the double dislocation where the phase jump adds up to $\pi$, and the amplitude at the CDW ordering wave vector should vanish by an identical argument to that of the half-vortex. We could again perform the FT to verify these results explicitly or, alternatively, argue that it should hold by way of the amount of dislocation charge associated with a given defect.
@@
We finish this section by analyzing cuts of the real space patterns of the double dislocation and the half-vortex across the vortex cores, seen in figure 7. The split peaks have a clear signature in the real space patterns. Since they arise due to jumps in phase of $\theta_{-}$ across the vortex core, the CDW pattern in real space showcases these phase jumps. To demonstrate this we place a waveform of the defect-free PDW in the background of the half-vortex and the double dislocation (dotted line).
@@
Starting with the double dislocation we can see the $\pi$ phase shift which occurs across the vortex core. The associated CDW pattern of the double dislocation has a “sawtooth” pattern, seen in figure 7b, to the left and to the right of the vortex core, which is even in $x$ being inversion symmetric. Comparison with the daughter $2Q$-CDW of the defect-free PDW gives us a subtle indication of the $\pi$ phase shift. A given sawtooth pattern lies within one of the waveforms of the defect-free PDW. Sufficiently far from the core of the vortex the teeth of the saws are odd in respect to the underlying waveform, peaking on the right side of the wave on the LHS and vice versa for the RHS of the vortex. This is indicative of a $\pi$ phase shift because the locations of the maximums and minimums of the double dislocation’s CDW change their relative orientation within the square wave, and is indeed needed to maintain inversion symmetry along the $x$-direction.
@@
Instead, for the half-vortex (figure 7a) the induced CDW on the LHS of the vortex is (basically) in phase with the background CDW. Again it is a sawtooth pattern, but the minima and maxima of the two waves coincide. Within the core of the vortex the two patterns slightly dephase from each other, and the half-vortex’s CDW is zero when the defect-free PDW amplitude is at a maximum, meaning there was a $\pi/2$ phase shift. This was to be expected from the above discussion. Unlike in the case of the double dislocation the half-vortex’s CDW is asymmetric about $x=0$. This too is to be expected from the breaking of inversion symmetry.
@@ 4.4 Tunneling DOS Spectra of the Topological Defects
We now compare and contrast experimental signatures associated with $dI/dV$ curves belonging to various superconducting order parameters. These curves illustrate that the half-vortex can be thought of as an interpolation between an FF-like state (in the core) to a LO-like state (at infinity). Also, evidence of inversion symmetry breaking in the tunneling data for the half-vortex state is discussed. We also changed the energy resolution in this section to .01.01.01.01 eV to smooth out the curves and make the low bias particle-hole symmetry more apparent.
@@
We begin with a comparison plot between the inequivalent sites for a defect-free LO state shown in figure 8. These plots showcase coherence peaks, like a uniform superconductor, low bias particle-hole symmetry as well as additional satellite peaks and a zero-bias electron DOS. Additionally, the electron tunneling data for the PDW state is periodic with half the period of the PDW since it senses the associated CDW (for detailed analysis see Ref. [56]), which is established by varying the $x$ coordinate. Thus the curves have a periodicity of four lattice spacings here, and not eight like our PDW order parameter, so we label the four representative curves with letters $A-D$. Alternatively, the periodicity is 4, and not 8, because the remaining four lattice sites have a SC gap which is $\pi$-phase shifted in respect to the first four, and a normal metal STM tip is blind to this effect. The full periodicity of the PDW state can be detected with pair (Jospehson) tunneling. This will be discussed elsewhere. The defect-free FF-state (not shown) possesses a constant tunneling-DOS across the entire plane, so only one representative curve is needed. Clearly then the FF state possesses data which is rotationally invariant (by $\pi/2$ in the CuO2 planes) unlike the above LO state.
@@
We now compare the LO states to a set of curves corresponding to the half-vortex (figure 9). The labels on each curve indicates how far we are from the vortex core in fractions of the halo radius, $r_{0}=24a_{0}$. We first note that near the core of the half-vortex the $\frac{dI}{dV}$ curve is not that of a free particle. In fact, it resembles a squeezed in FF-state, even possessing the discrete rotational symmetry of the lattice (not shown). As we travel out to the edge of the vortex, the half-vortex begins to become more LO-like than FF-like. This is the interpolation from a FF to a LO state that the half-vortex undergoes, apparent from our boundary conditions in equation (3.16).
@@
The presence of a half-flux quanta will further distinguish the half-vortex from the defect-free PDW far outside the vortex core. In appendix E we provide supplementary $dI/dV$ plots taken outside the core of the vortex which indicated there is a relative shift in the tunneling DOS curves. Indeed, figure 14 compares the half-vortex tunneling spectrum to the right and to the left of the vortex core (figures 14a and 14b, respectively). Again there is a periodicity present (outside the vortex core) in these plots, and just like an LO state: the $dI/dV$ curves repeat every four lattice spacings. It should also be noted that the curves presented in the appendix have a defect which is placed on the CuO2 bonds opposed to at the center of the plaquette. This changes the appearance of the $dI/dV$ curves, but the spectral weight associated with the quasi-particles remains the same. That is, we shifted the period $4$ CDW by half a lattice spacing, which costs us no, or very little, energy to do so.
@@
The half-vortex becomes more LO-like outside the core of the half-vortex, but since this topological defect breaks inversion symmetry, it’s tunneling DOS must reflect this, unlike the LO state. A smoking gun signature of inversion symmetry breaking is present by comparing the $dI/dV$ curves to the left and to the right of the vortex core (figures 14a and 14b, respectively). It is apparent the data on the left is $2$ lattice spacings behind the right (or vice versa), which we attribute to the jump in $\theta_{+}$ by $\pi/2$ across the vortex core. Indeed, the accumulated phase belonging to a Cooper pair with nonzero COM-momentum: $\mathbf{Q}\cdot\mathbf{r}=\pi/2$ if $\mathbf{r}=2a_{0}\mathbf{e}_{x}$. We will see another example of inversion symmetry breaking in the next section when we discuss the spectral functions for the half-vortex. This jump in phase distinguishes the half-vortex from a defect-free LO state, even outside the half-vortex core.
@@
Note the full-vortex also possesses a phase jump of $\Delta\theta_{+}=\pi$ across the vortex core, but this gives a relative shift of four lattice spacings when comparing electron tunneling DOS on the left and right hand sides of the full-vortex. This means there is no analogous signature belonging to the full-vortex as there was for the half-vortex when using a normal metal tip. This is simply because the LO state has the same periodicity as this shift. On the other hand, in the case of a superconducting tip there will be a difference in the pair tunneling DOS on the right and left hand sides of the full-vortex. Finally, we note that the presence of a gap in the core of the half-vortex distinguishes it from the other two topological defects since the latter two have a vanishing gap here. This results in a free particle tunneling DOS in the core of the defect for the double dislocation and the full-vortex opposed to a squeezed in FF state seen in the core of the half-vortex (see figure 15a). Finally, we can see in figure 15b that the half-vortex possesses additional satellites peaks past these other two defects, and so the full-vortex and the double dislocation more closely resemble the pure PDW outside the core. The inequivalent gaps corresponding to the distinct Fourier components of the half-vortex is responsible for these additional satellites.

### v4

@@ 4.3 Patterns of the FT of the topological defects
We now analyze in detail the effects that the half-vortex and the double dislocation have on the induced CDW order. Recall that inside the core of the half-vortex the SC order parameter is mostly FF type, since the amplitude of one of the two PDW order parameters must vanish at the location of the half-vortex, i.e., the origin (see Sec. 3). Since the FF state breaks inversion symmetry in the $x$-direction, the corresponding CDW pattern inherits this broken symmetry. Also, the edge dislocations of the CDW order parameter break inversion symmetry in the $y$-direction. Note that, the parity operator, in the $x$-direction, changes the location of the branch cuts and the signs of the winding numbers when it acts on the order parameters, which changes the sign of the dislocation charge. As a result, the corresponding charge density patterns are flipped on their head under this operation. Since the location of the branch cuts have no physically significant effects on the charge density, the full vortex is invariant under this operation.
The ordering wave vector $2\mathbf{Q}$ has many features made more apparent in the Fourier transforms of the LDOS, which are shown in Figs. 6a and 6b. Notice that the Fourier transforms of the LDOS for half-vortex and the double dislocation feature split peaks at the $2\mathbf{Q}$ ordering wave vector where the amplitude of the FT-LDOS is zero there. In contrast, in the case of the Abrikosov vortex, the Fourier transform of the LDOS is just the transform of $\cos(2\mathbf{Q}x)$ which has a single peak at the ordering wave vector. Cuts of the FT-LDOS along $k_{y}=0$ are given in Figs. 6c and 6d to more clearly illustrate these split peaks.
@@
The split peaks are signature of the defects of the PDW phase associated with jumps in the phase $\theta_{-}$, defined in Eq.  (2.8), across the core of the topological defect. For example, across the half-vortex the phase jumps by ±$\pm\pi/2$ since it winds by $\pi$ around the half-vortex. This implies that the $2Q$ Fourier component is equal to itself times $i$ across the core of the defect, suggesting that this Fourier component must be zero. We can explicitly verify this prediction by examining the FT of $\rho_{2\mathbf{Q}}(\mathbf{r})$. It is apparent that a nonzero CDW winding number is responsible for the vanishing of $\rho_{2\mathbf{Q}}(\mathbf{r})$ at the center of the defect. In other words, the phase shift that causes this destructive interference is a measurement of the Burgers vector, which is the topological charge of the dislocation. Similar interference patterns are well known to exist in electron diffraction in crystals of semiconductors with dislocations.
@@
Similar split peaks in the Fourier transform of the tunneling LDOS were also predicted to exist at the PDW halo of an Abrikosov vortex of a superconductor in Refs. [16, 64], but their physical origin is very different. Indeed, in the case of the vortex halo there is a phase shift in the $\rho_{\mathbf{Q}}(\mathbf{r})$ (instead of $\rho_{2\mathbf{Q}}(\mathbf{r})$) Fourier component of the local charge density caused the Abrikosov vortex of the uniform component of the superconductor.
@@
Figure 8: Comparison of $\frac{dI}{dV}$ curves related to the electron tunneling DOS for a period $8$ PDW (LO state). The electron tunneling DOS is sensitive to the periodicity of the associated period $4$ CDW. The labeled $A-D$ are the tunneling DOS traces at the four inequivalent sites of the CDW. Each consecutive curve is offset by 1 for clarity. Notice that in the PDW traces particle-hole symmetry is present only at low bias (low energies).
@@
Figure 9: A set of $\frac{dI}{dV}$ curves for the half-vortex belonging to different lattice sites. The labels indicate the distance the lattice site in question is from the vortex core in fractions of the halo radius, $r_{0}=24a_{0}$. The top two curves are shifted up by $0.75$ and $1.5$ units in respect to zero. Close to the center of the core the curves resemble a squeezed in FF state more so than an LO state, but as we move outward the coherence peaks grow, and the half-vortex behaves more like an LO state. In appendix E, Fig. 14 additional curves corresponding to a half-vortex are provided which indicates a 4 lattice site periodicity, like the LO state, but there is also a shift in tunneling spectra associated with the jump in phase of $\theta_{+}$ across the vortex core. This, along with reduced coherence peaks and additional satellites, distinguished the half-vortex from the pure PDW even far outside the vortex core.
@@
Figure 10: The Bogoliubov Fermi surface, $A(\mathbf{k},\omega=0)$, for a PDW state with a half-vortex of radius $r_{0}=24a_{0}$ and a $d$-wave form factor. In (a) we plot the upper portion of the Bogoliubov Fermi surface, which indicates a redistribution of spectral weight in respect to the defect-free PDW (see Appendix F for plots). In (b) we zoom in on these loops, which indicate inversion symmetry breaking. We also partition the Fermi surface of the half-vortex into (c) the particlelike portion and (d) the holelike portion. Notice the intensity of the spectral weight belonging to the Fermi arcs tend to be either particlelike or holelike depending on the side you are on.
@@
Alternatively, when there is a winding number in both PDW components, we can picture the phase jump as occurring in both winding numbers independently. Recalling the form of the induced $2Q$-CDW from Eq.  (2.6), we see that complex conjugation flips the phase-winding of one of the PDW components. Thus, when we have a vortex in both components $\Delta_{\pm\bm{Q}}$ with the same winding, i.e., the Abrikosov vortex, the phase differences cancel each other, and there is no net phase jump across the core of the vortex. Hence, in the case of the Abrikosov vortex there is not a split peak. Equivalently, the Abrikosov vortex does not have any dislocation charge associated with it to cause a split peak to exist. We can contrast this with the double dislocation where the phase jump adds up to $\pi$, and the amplitude at the CDW ordering wave vector should vanish by an identical argument to that of the half-vortex. We could again perform the FT to verify these results explicitly or, alternatively, argue that it should hold by way of the amount of dislocation charge associated with a given defect.
@@
We finish this section by analyzing cuts of the real space patterns of the double dislocation and the half-vortex across the vortex cores, seen in Fig. 7. The split peaks have a clear signature in the real space patterns. Since they arise due to jumps in phase of $\theta_{-}$ across the vortex core, the CDW pattern in real space showcases these phase jumps. To demonstrate this we place a waveform of the defect-free PDW in the background of the half-vortex and the double dislocation (dotted line).
@@
Starting with the double dislocation we can see the $\pi$ phase shift which occurs across the vortex core. The associated CDW pattern of the double dislocation has a “sawtooth” pattern, seen in Fig. 7b, to the left and to the right of the vortex core, which is even in $x$ being inversion symmetric. Comparison with the daughter $2Q$-CDW of the defect-free PDW gives us a subtle indication of the $\pi$ phase shift. A given sawtooth pattern lies within one of the waveforms of the defect-free PDW. Sufficiently far from the core of the vortex the teeth of the saws are odd in respect to the underlying waveform, peaking on the right side of the wave on the l.h.s. and vice versa for the r.h.s. of the vortex. This is indicative of a $\pi$ phase shift because the locations of the maximums and minimums of the double dislocation’s CDW change their relative orientation within the square wave, and is indeed needed to maintain inversion symmetry along the $x$-direction.
@@
Instead, for the half-vortex [Fig. 7a] the induced CDW on the l.h.s. of the vortex is (basically) in phase with the background CDW. Again it is a sawtooth pattern, but the minima and maxima of the two waves coincide. Within the core of the vortex the two patterns slightly dephase from each other, and the half-vortex’s CDW is zero when the defect-free PDW amplitude is at a maximum, meaning there was a $\pi/2$ phase shift. This was to be expected from the above discussion. Unlike in the case of the double dislocation the half-vortex’s CDW is asymmetric about $x=0$. This too is to be expected from the breaking of inversion symmetry.
@@ 4.4 Tunneling DOS Spectra of the Topological Defects
We now compare and contrast experimental signatures associated with $dI/dV$ curves belonging to various superconducting order parameters. These curves illustrate that the half-vortex can be thought of as an interpolation between an FF-like state (in the core) to a LO-like state (at infinity). Also, evidence of inversion symmetry breaking in the tunneling data for the half-vortex state is discussed. We also changed the energy resolution in this section to $0.01$ eV to smooth out the curves and make the low bias particle-hole symmetry more apparent.
@@
We begin with a comparison plot between the inequivalent sites for a defect-free LO state shown in Fig. 8. These plots showcase coherence peaks, like a uniform superconductor or low-bias particle-hole symmetry as well as additional satellite peaks and a zero-bias electron DOS. Additionally, the electron tunneling data for the PDW state is periodic with half the period of the PDW since the data sense the associated CDW (for a detailed analysis see Ref. [55]), which is established by varying the $x$ coordinate. Thus the curves have a periodicity of four lattice spacings here, and not eight like our PDW order parameter, so we label the four representative curves with letters $A-D$. Alternatively, the periodicity is 4, and not 8, because the remaining four lattice sites have a SC gap which is $\pi$-phase shifted in respect to the first four, and a normal metal STM tip is blind to this effect. The full periodicity of the PDW state can be detected with pair (Jospehson) tunneling. This will be discussed elsewhere. The defect-free FF state (not shown) possesses a constant tunneling-DOS across the entire plane, so only one representative curve is needed. Clearly then the FF state possesses data which are rotationally invariant (by $\pi/2$ in the CuO2 planes) unlike the above LO state.
@@
We now compare the LO states to a set of curves corresponding to the half-vortex (Fig. 9). The labels on each curve indicates how far we are from the vortex core in fractions of the halo radius, $r_{0}=24a_{0}$. We first note that near the core of the half-vortex the $\frac{dI}{dV}$ curve is not that of a free particle. In fact, it resembles a squeezed in FF state, even possessing the discrete rotational symmetry of the lattice (not shown). As we travel out to the edge of the vortex, the half-vortex begins to become more LO-like than FF-like. This is the interpolation from a FF to a LO state that the half-vortex undergoes, apparent from our boundary conditions in Eq.  (3.16).
@@
The presence of a half-flux quanta will further distinguish the half-vortex from the defect-free PDW far outside the vortex core. In Appendix E we provide supplementary $dI/dV$ plots taken outside the core of the vortex, which indicated there is a relative shift in the tunneling DOS curves. Indeed, Fig. 14 in Appendix E compares the half-vortex tunneling spectrum to the right and to the left of the vortex core [Figs. 14a and 14b, respectively]. Again there is a periodicity present (outside the vortex core) in these plots, and just like an LO state, the $dI/dV$ curves repeat every four lattice spacings. It should also be noted that the curves presented in the appendix have a defect which is placed on the CuO2 bonds opposed to at the center of the plaquette. This changes the appearance of the $dI/dV$ curves, but the spectral weight associated with the quasiparticles remains the same. That is, we shifted the period $4$ CDW by half a lattice spacing, which costs us no, or very little, energy to do so.
@@
The half-vortex becomes more LO-like outside the core of the half-vortex, but since this topological defect breaks inversion symmetry, its tunneling DOS must reflect this, unlike the LO state. A smoking gun signature of inversion symmetry breaking is present by comparing the $dI/dV$ curves to the left and to the right of the vortex core [figures 14a and 14b, respectively]. It is apparent the data on the left are two lattice spacings behind the right (or vice versa), which we attribute to the jump in $\theta_{+}$ by $\pi/2$ across the vortex core. Indeed, the accumulated phase belonging to a Cooper pair with nonzero COM momentum: $\mathbf{Q}\cdot\mathbf{r}=\pi/2$ if $\mathbf{r}=2a_{0}\mathbf{e}_{x}$. We will see another example of inversion symmetry breaking in the next section when we discuss the spectral functions for the half-vortex. This jump in phase distinguishes the half-vortex from a defect-free LO state, even outside the half-vortex core.
@@
Note the full vortex also possesses a phase jump of $\Delta\theta_{+}=\pi$ across the vortex core, but this gives a relative shift of four lattice spacings when comparing electron tunneling DOS on the left and right hand sides of the full vortex. This means there is no analogous signature belonging to the full vortex as there was for the half-vortex when using a normal metal tip. This is simply because the LO state has the same periodicity as this shift. On the other hand, in the case of a superconducting tip there will be a difference in the pair tunneling DOS on the right- and left-hand sides of the full vortex. Finally, we note that the presence of a gap in the core of the half-vortex distinguishes it from the other two topological defects since the latter two have a vanishing gap here. This results in a free particle tunneling DOS in the core of the defect for the double dislocation and the full vortex opposed to a squeezed in FF state seen in the core of the half-vortex [see Fig. 15a in Appendix E]. Finally, we can see in Fig. 15b that the half-vortex possesses additional satellites peaks past these other two defects, and so the full vortex and the double dislocation more closely resemble the pure PDW outside the core. The inequivalent gaps corresponding to the distinct Fourier components of the half-vortex is responsible for these additional satellites.